Grade 10Engineering Entrance Exams

Sin2x. + Cos2y is equal to =. If x+y =90° and x=2y:

Gourav sharma
9 Years agoGrade 10

1 Answer

Robo
9 Years ago
Yo ...Solve x+y=π/2 and x=2yYou will get x= π/3 and y= π/6 Put them in sin2x + cos2y.Hope this hepls

question mark
Hello,
 
Can we learn and prepare for MHT-CET online ?
engineering entrance exams

Last Activity: 4 Years ago